# TopoBDA: Towards Bezier Deformable Attention for Road Topology Understanding

### v2

@@ 2.3.2 Road Topology
TopoNet [27] proposes a graph neural network architecture that models the relationships between centerlines and traffic elements. It incorporates prior relational knowledge to enhance feature interactions. TopoMLP [66] introduces an advanced pipeline for understanding driving topology by incorporating lane coordinates into the topology framework and using an L1 loss function to refine the interaction points. CGNet [14] focuses on preserving the continuity of centerline graphs and improving topology accuracy through modules like Junction Aware Query Enhancement and Bezier Space Connection. Topo2D [25] integrates 2D lane priors to improve 3D lane detection and topology reasoning. TopoLogic [12] proposes managing lane topology relationships by combining geometric lane distance with similarity-based topology relationships. TopoFormer [47] introduces a lane aggregation layer that leverages geometric distance in driving self-attention, along with a counterfactual intervention layer to improve reasoning by considering alternative scenarios and their causal impacts.
@@ 3 Methodology
TopoBDA (Figure 2) starts by extracting Bird’s Eye View (BEV) features from multi-camera 360-degree imagery. A transformer decoder then processes these BEV features within its cross-attention mechanism using a sparse query approach. In this approach, each query corresponds to a centerline instance rather than individual points, which improves computational efficiency. Subsequently, the decoder predicts Bezier control points for each instance query which are then converted into dense polyline points via matrix multiplication. The Bezier representation offers a compact and compressed formulation, reducing computational complexity at the regression heads [11, 53, 27]. Additionally, each centerline query predicts an instance mask for auxiliary loss, though these masks are not utilized during inference.
@@
In the TopoBDA architecture, the BEV feature extraction (Figure 2) starts with a set of $N$ images, $\{\mathbf{I}_{i}\}_{i=1}^{N}$, each having a dimension of $H_{I}\times W_{I}\times 3$, with $H_{I}$ and $W_{I}$ representing the height and width, respectively. These images are first converted into perspective view features, $\{\mathbf{F}_{PV_{i}}\}_{i=1}^{N}$, using a feature extraction function $f_{PV}$, such that $\mathbf{F}_{PV_{i}}=f_{PV}(\mathbf{I}_{i})$ and $\mathbf{F}_{PV_{i}}\in\mathbb{R}^{H_{PV}\times W_{PV}\times C_{PV}}$. These perspective view features are then aggregated and projected into a single Bird’s Eye View (BEV) feature map, $\mathbf{F}_{BEV}$, using a projection function $f_{BEV}$. This function can be implemented via Lift-Splat-Shoot (LSS) [51, 18, 30], transformers [31, 6, 81, 62], or other projection techniques [29, 67, 15]. This projection is defined as $\mathbf{F}_{BEV}=f_{BEV}(\{\mathbf{F}_{PV_{i}}\}_{i=1}^{N})$, where $\mathbf{F}_{BEV}\in\mathbb{R}^{H_{BEV}\times W_{BEV}\times C_{BEV}}$. This methodology effectively captures and transforms spatial features into the BEV space, facilitating further analysis and processing within the TopoBDA framework.
Next, the extracted BEV features $\mathbf{F}_{BEV}$ are fed into the transformer decoder, where they are used as the value input in the cross-attention mechanism. Each query, corresponding to a single centerline instance, interacts with $\mathbf{F}_{BEV}$, and Bezier control points are predicted at each decoder layer. The attention mechanism in TopoBDA is based on Bezier Deformable Attention (BDA) and it is guided by these predicted Bezier control points. Further details of this mechanism are provided in Section 3.1 and the overall structure of the TopoBDA transformer decoder is provided in Section 3.2.
@@ 3.1 Towards Bezier Deformable Attention
This section explains the transition from Single-Point Deformable Attention (SPDA) to Multi-Point Deformable Attention (MPDA) for methodologies based on Bezier keypoint regression (or Bezier control point regression). This transition replaces traditional deformable attention heads with dense polyline points obtained via matrix multiplication on Bezier control points (Figure 3a).
Next, the concept of Bezier Deformable Attention (BDA) is introduced. BDA enhances the deformable attention mechanism by using Bezier control points. In this approach, traditional attention heads are substituted with the control points of a polyline curve, facilitating more flexible and adaptive attention. In each layer, the predicted control points are utilized to guide the Bezier Deformable Attention process (Figure 3b).
3.1.1 Bezier Curve Representation
@@
3.1.2 Deformable Attention Mechanism
@@ 3.1.3 Adaptation of Bezier Regression Methods to Multi-Point Deformable Attention (MPDA)
Figure 3: Comparison of Multi-Point Deformable Attention (MPDA) and Bezier Deformable Attention (BDA): MPDA necessitates an additional matrix multiplication block within each transformer decoder. Despite their different input utilizations, the mechanisms of MPDA and BDA are fundamentally the same.
This methodology involves extracting $L+1$ polyline points $\mathbf{P}=\{\mathbf{p}_{0},\mathbf{p}_{1},\ldots,\mathbf{p}_{L}\}$ from the control points $\mathbf{C}=\{\mathbf{c}_{0},\mathbf{c}_{1},\ldots,\mathbf{c}_{N}\}$ through Bezier extraction process. These polyline points are then utilized in the deformable attention mechanism. This approach replaces the traditional multi-head attention mechanism, which operates on a single reference point, with a more flexible and adaptive mechanism. Instead of using multiple heads $M$ on a single point $\mathbf{p}$, this method employs a single head for each of the dense polyline points $\mathbf{p}_{l}$. By utilizing multiple points distributed along the polyline, this technique captures the thin, elongated characteristics of polylines, thereby enhancing the attention mechanism’s ability to model complex patterns and dependencies.
@@
The polyline points $\mathbf{p}_{l}$ can be extracted from the control points using the following formula, which is derived from Eq. (1) and Eq. (2):
@@
for $l=0,1,\ldots,L$, where $t_{l}$ are uniformly spaced within the interval $[0,1]$. Eq. (5) demonstrates that adapting MPDA to the Bezier keypoint concept necessitates converting Bezier control points into polyline points. This conversion is implemented through matrix multiplication in each decoder layer, as illustrated in Figure 3a. The detailed process of this implementation is provided in Section S.1 of the supplementary materials.
@@ 3.1.4 Bezier Deformable Attention
BDA enhances the traditional multi-head attention mechanism by replacing the single-point focus $\mathbf{p}$ with control points $\mathbf{c}_{n}$ of the Bezier curve. Unlike MPDA, which relies on dense polyline points, BDA directly uses these control points as reference points within the deformable attention mechanism. This approach eliminates the need for converting control points to polyline points in each decoder layer (Figure 3), reducing computational complexity. By predicting Bezier control points and focusing attention around them, BDA improves the learning process, leading to more effective and accurate predictions.
@@ 3.2 Bezier Deformable Attention Based Transformer Decoder
The overview of the general pipeline of each layer in the transformer decoder of TopoBDA is shown in Figure 4. The process involves several steps:
@@ 3.3.1 Instance Mask Formulation as an Auxiliary Loss
To achieve this, centerlines are converted into masks using a parameter $W$ in the BEV domain. Utilizing the generated ground truth instance masks, each centerline query in the transformer decoder predicts both Bezier control points and a mask probability map for each centerline instance (Figure 5).
@@ 3.3.2 Mask-L1 Mix Matcher
When $\lambda_{\text{reg}}=0$, it operates as a mask matcher, whereas when $\lambda_{\text{mask}}=0$, it operates as an L1 matcher. When both $\lambda_{\text{reg}}$ and $\lambda_{\text{mask}}$ are non-zero, it is referred to as a Mask-L1 mix matcher. This hybrid matching approach, inspired by the Mask DINO framework [24], is adapted to enhance road topology understanding. Experimental results show that utilizing the instance mask formulation in both the main loss and bipartite matching (Hungarian matcher) improves performance. Furthermore, the Mask-L1 mix matcher accelerates the convergence of topology performance.
@@ 3.4 Fusion Methodology
Sensor fusion for road topology understanding has not been extensively explored in the literature. In this work, the raw features of radar and lidar point clouds are directly concatenated at the voxel space level. While similar to the BEVFusion studies [32, 41, 58] it differs by performing fusion in voxel space instead of fusing in BEV space to preserve more information. In contrast, SDMap fusion is done in the BEV space, and SDMap features are concatenated with BEV features from other sensors. The mathematical background of the proposed multi-sensor fusion pipeline is detailed in Section S.2 of the supplementary materials.
@@ 3.5 Auxiliary One-to-Many Set Prediction Loss Strategy
To improve training efficiency, an auxiliary one-to-many set prediction loss strategy [20, 35] is employed. This study, to the best of our knowledge, is the first to perform ablation experiments on this strategy for road topology understanding. Results show that this approach significantly enhances the architecture’s ability to understand road topology.
The auxiliary one-to-many set prediction loss approach involves a smaller decoder and a larger decoder with shared weights. The smaller decoder uses the ground truth set directly, while the larger decoder employs a repeated ground truth set to increase the number of positive samples. During training, both decoders are utilized; while only the smaller decoder is employed during inference. The mathematical background of this loss strategy is provided in Section S.3 of the supplementary materials.
@@ 4 Experimental Evaluation
This section provides a comprehensive analysis of the methodologies employed in this study. It commences with a detailed description of the datasets and metrics (Section 4.1). Subsequently, Section 4.2 presents an extensive experimental evaluation, encompassing an analysis of instance mask formulation, comparisons of various attention mechanisms, an examination of auxiliary one-to-many set prediction loss, an efficiency analysis of the proposed Bezier deformable attention, and a study on multi-modal fusion involving camera, radar, lidar, and SDMap. The results are juxtaposed with state-of-the-art methods on both OpenLane-V1 and OpenLane-V2.
Supplementary material provides additional details on the various loss functions (Section S.4), the implementation specifics of the proposed architecture (Section S.5), and further experiments (Section S.6) including view transformation, backbone variations, number of epochs, number of encoder and decoder layers, and efficient multi-scale implementation ablations.
@@ 4.1.1 Datasets
Figure 6: The PV and BEV samples from the OpenLane-V2 dataset. The top row and bottom row shows an example from Subset-A and Subset-B, respectively. (a) show the PV, highlighting centerlines and traffic elements. (b) displays the corresponding centerline instances in the BEV domain, and (c) displays centerlines with colors highlighting the topological relationships between centerlines and traffic elements. (d) demonstrate the topological relationships between different centerlines.
The OpenLane-V2 dataset [59] is a centerline detection and road topology understanding dataset and has two subsets: Subset-A and Subset-B. Subset-A is derived from the Argoverse 2 (AV2) [65] dataset, containing samples from six cities (e.g., Miami, Pittsburgh, Austin), collected using a seven-camera setup with the front camera positioned vertically. This subset contains 22,477 training samples, 4,806 validation samples, and 4,816 test samples, all with a resolution of $2048\times 1550$ pixels. The top row of Figures 6 shows the PV and BEV images of a random sample in Subset-A. The front view of the PV is cropped due to the vertical camera positioning. Multi-modal fusion studies are conducted with lidar data and SDMap and no radar data is available in Subset-A.
Subset-B is based on the NuScenes [2] dataset and contains data collected from Boston and Singapore, using a six-camera setup. It includes 27,968 training, 6,019 validation, and 6,008 test samples, with an image resolution of $1600\times 900$ pixels. Subset-B has a higher proportion of night and rainy scenes compared to Subset-A but lacks ground height information. Sensor fusion studies are implemented with both radar and lidar data but no available SDMap information in this subset. The bottom row of Figures 6 displays the PV and BEV images of a random sample from Subset-B.
The OpenLane-V1 dataset [5] is a comprehensive benchmark for 3D lane detection, derived from the Waymo Open dataset. It consists of 1000 segments with 200K frames, captured under diverse conditions at $1920\times 1280$ resolution. This dataset provides diverse and challenging scenarios for evaluating lane detection algorithms, including a wide range of weather conditions, lighting variations, and road types.

### v3

@@ 2.3.2 Road Topology
TopoNet [li_graph-based_2023] proposes a graph neural network architecture that models the relationships between centerlines and traffic elements. It incorporates prior relational knowledge to enhance feature interactions. TopoMLP [wu_topomlp_2024] introduces an advanced pipeline for understanding driving topology by incorporating lane coordinates into the topology framework and using an L1 loss function to refine the interaction points. CGNet [han_continuity_2024] focuses on preserving the continuity of centerline graphs and improving topology accuracy through modules like Junction Aware Query Enhancement and Bezier Space Connection. Topo2D [li_enhancing_2024] integrates 2D lane priors to improve 3D lane detection and topology reasoning. TopoLogic [fu_topologic_2024] proposes managing lane topology relationships by combining geometric lane distance with similarity-based topology relationships. TopoFormer [lv_t2sg_2024] introduces a lane aggregation layer that leverages geometric distance in driving self-attention, along with a counterfactual intervention layer to improve reasoning by considering alternative scenarios and their causal impacts.
@@ 3 Methodology
TopoBDA (Figure 2) starts by extracting Bird’s Eye View (BEV) features from multi-camera 360-degree imagery. A transformer decoder then processes these BEV features within its cross-attention mechanism using a sparse query approach. In this approach, each query corresponds to a centerline instance rather than individual points, which improves computational efficiency. Subsequently, the decoder predicts Bezier control points for each instance query, which are then converted into dense polyline points via matrix multiplication. The Bezier representation offers a compact and compressed formulation, reducing computational complexity at the regression heads [feng_rethinking_2022, qiao_end--end_2023, li_graph-based_2023]. Additionally, each centerline query predicts an instance mask for auxiliary loss, though these masks are not utilized during inference.
@@
In the TopoBDA architecture, the BEV feature extraction (Figure 2) starts with a set of $N$ images, $\{\mathbf{I}_{i}\}_{i=1}^{N}$, each having a dimension of $H_{I}\times W_{I}\times 3$, with $H_{I}$ and $W_{I}$ representing the height and width, respectively. These images are first converted into perspective view features, $\{\mathbf{F}_{PV_{i}}\}_{i=1}^{N}$, using a feature extraction function $f_{PV}$, such that $\mathbf{F}_{PV_{i}}=f_{PV}(\mathbf{I}_{i})$ and $\mathbf{F}_{PV_{i}}\in\mathbb{R}^{H_{PV}\times W_{PV}\times C_{PV}}$. These perspective view features are then aggregated and projected into a single Bird’s Eye View (BEV) feature map, $\mathbf{F}_{BEV}$, using a projection function $f_{BEV}$. This function can be implemented via Lift-Splat-Shoot (LSS) [philion_lift_2020, huang_bevdet_2021, li_bevdepth_2023], transformers [li_bevformer_2024, chen_efficient_2022, zhou_cross-view_2022, wang_exploring_2023], or other projection techniques [li_fast-bev_2024, xie_m2bev_2022, harley_simple-bev_2023]. This projection is defined as $\mathbf{F}_{BEV}=f_{BEV}(\{\mathbf{F}_{PV_{i}}\}_{i=1}^{N})$, where $\mathbf{F}_{BEV}\in\mathbb{R}^{H_{BEV}\times W_{BEV}\times C_{BEV}}$. This methodology effectively captures and transforms spatial features into the BEV space, facilitating further analysis and processing within the TopoBDA framework.
Next, the extracted BEV features $\mathbf{F}_{BEV}$ are fed into the transformer decoder, where they are used as the value input in the cross-attention mechanism. Each query, corresponding to a single centerline instance, interacts with $\mathbf{F}_{BEV}$, and Bezier control points are predicted at each decoder layer. The attention mechanism in TopoBDA is based on Bezier Deformable Attention (BDA), and it is guided by these predicted Bezier control points. Further details of this mechanism are provided in Section 3.1, and the overall structure of the TopoBDA transformer decoder is provided in Section 3.2.
@@ 3.1 Towards Bezier Deformable Attention
Figure 3: Comparison of Single-Point Deformable Attention (SPDA), Multi-Point Deformable Attention (MPDA), and Bezier Deformable Attention (BDA). Points denote the reference positions (anchors) for each attention head, while arrows indicate the learned offsets that shift attention from these anchors to the actual sampling locations where features are aggregated. SPDA uses identical reference positions across all heads, whereas MPDA and BDA employ distinct reference positions per head, improving attention efficiency for polyline structures. Although MPDA and BDA share the same underlying mechanism, they differ in how multiple reference points $(p_{x},p_{y})$ are selected. BDA directly utilizes Bezier points as reference positions, while MPDA requires conversion of Bezier points into polyline points and utilizes polyline points as reference positions.
This section describes the transition from Single-Point Deformable Attention (SPDA) to Multi-Point Deformable Attention (MPDA) and Bezier Deformable Attention (BDA) in methodologies based on Bezier keypoint regression. Unlike SPDA, which relies on a single reference point per attention head, MPDA and BDA introduce multiple reference points, enabling more expressive and adaptive feature aggregation for complex polyline structures (Figure 3).
Although MPDA and BDA share the same underlying mechanism, they differ in how reference points are selected. MPDA requires conversion of Bezier control points into dense polyline points via matrix multiplication, which are then used as reference positions for attention heads (Figure 4a). In contrast, BDA directly utilizes Bezier control points as reference positions, eliminating the need for conversion and preserving the original regression targets predicted by the model (Figure 4b). In each layer, the predicted control points guide the Bezier Deformable Attention process.
The following subsections provide an overview of Bezier curve representation and the standard Single-Point Deformable Attention (SPDA), followed by the adaptation of Multi-Point Deformable Attention (MPDA) to Bezier control points and the introduction of Bezier Deformable Attention (BDA).
3.1.1 Recap of Bezier Curve Representation
@@
3.1.2 Recap of Deformable Attention Mechanism
@@ 3.1.3 Adaptation of Bezier Regression Methods to Multi-Point Deformable Attention (MPDA)
MPDA replaces the traditional multi-head attention mechanism (SPDA), which operates on a single reference point, with a more flexible and adaptive mechanism. Instead of using multiple heads $M$ on a single point $\mathbf{p}$, this method employs a single head for each of the dense polyline points $\mathbf{p}_{l}$ as shown in Figure 3. By utilizing multiple points distributed along the polyline, this technique captures the thin, elongated characteristics of polylines, thereby enhancing the attention mechanism’s ability to model complex patterns and dependencies.
@@
This methodology necessitates extracting $L+1$ polyline points $\mathbf{P}=\{\mathbf{p}_{0},\mathbf{p}_{1},\ldots,\mathbf{p}_{L}\}$ from the control points $\mathbf{C}=\{\mathbf{c}_{0},\mathbf{c}_{1},\ldots,\mathbf{c}_{N}\}$ through Bezier extraction process. The polyline points $\mathbf{p}_{l}$ can be extracted from the control points using the following formula, which is derived from Eq. (1) and Eq. (2):
@@
for $l=0,1,\ldots,L$, where $t_{l}$ are uniformly spaced within the interval $[0,1]$. This conversion is implemented through matrix multiplication in each decoder layer, as illustrated in Figure 4a. The detailed process of this implementation is provided in Supplementary Section S.1 of the supplementary materials.
Figure 4: Comparison of Multi-Point Deformable Attention (MPDA) and Bezier Deformable Attention (BDA): MPDA necessitates an additional matrix multiplication block within each transformer decoder to convert predicted Bezier control points into polyline points for use as reference points. Despite their different input utilizations as reference points, the mechanisms of MPDA and BDA blocks are fundamentally the same: each attention head operates on a distinct reference point—polyline points in MPDA and Bezier control points in BDA.
@@ 3.1.4 Bezier Deformable Attention
BDA enhances the traditional multi-head attention mechanism by replacing the single-point focus $\mathbf{p}$ with control points $\mathbf{c}_{n}$ of the Bezier curve. From this perspective, the inherent mechanism of BDA is the same with MPDA, as shown in Figure 3. Unlike MPDA, which relies on dense polyline points, BDA directly uses these control points as reference points within the deformable attention mechanism, eliminating the need for converting control points to polyline points in each decoder layer and reducing computational complexity slightly (Figure 4). By predicting Bezier control points and focusing attention around them, BDA improves the learning process, leading to more effective and accurate predictions.
@@
From a theoretical computational complexity standpoint, there is no difference between the attention mechanism of SPDA and BDA, as the only modification lies in the reference points that guide attention across different heads. However, SPDA requires learning an additional reference point learning mechanism in polyline detection structures, which increases the complexity slightly.
To summarize, the key distinction between SPDA, MPDA, and BDA lies in how reference points are utilized across attention heads. From a computational complexity perspective, these mechanisms are nearly equivalent when the number of attention heads is the same. However, SPDA introduces slight overhead due to its additional reference point learning, while MPDA incurs cost from converting control points into polyline points. BDA avoids this by directly using Bezier control points.
@@ 3.2 Bezier Deformable Attention Based Transformer Decoder
The overview of the general pipeline of each layer in the transformer decoder of TopoBDA is shown in Figure 5. The step-wise operation detail of the TopoBDA decoder is also shown in Supplementary Algorithm 1 in Section S.7. The iterative refinement-based Bezier control points prediction process involves several steps:
@@ 3.3.1 Instance Mask Formulation as an Auxiliary Loss
To achieve this, centerlines are converted into masks using a parameter $W$ in the BEV domain. Utilizing the generated ground truth instance masks, each centerline query in the transformer decoder predicts both Bezier control points and a mask probability map for each centerline instance (Figure 6). Supplementary Algorithm 1 in Section S.7 demonstrates the algorithmic details about the mask probability map generation within the TopoBDA transformer decoder.
@@ 3.3.2 Mask-L1 Mix Matcher
When $\lambda_{\text{reg}}=0$, it operates as a mask matcher, whereas when $\lambda_{\text{mask}}=0$, it operates as an L1 matcher. When both $\lambda_{\text{reg}}$ and $\lambda_{\text{mask}}$ are non-zero, it is referred to as a Mask-L1 mix matcher. This hybrid matching approach, inspired by the Mask DINO framework [li_mask_2023], is adapted to enhance road topology understanding. Experimental results show that utilizing the instance mask formulation in both the main loss and bipartite matching (Hungarian matcher) improves performance. Furthermore, the Mask-L1 mix matcher accelerates the convergence of topology performance.
@@ 3.4 Fusion Methodology
Figure 7: Sensor fusion pipeline in the TopoBDA architecture.
Sensor fusion for road topology understanding has not been extensively explored in the literature. The sensor and SDMap fusion pipeline used in TopoBDA is illustrated in Figure 7. In TopoBDA, camera and lidar features are first concatenated in the voxel space to preserve spatial granularity. This approach differs from BEVFusion studies [liang_bevfusion_2022, liu_bevfusion_2023, tang_multi-modality_2023], which typically perform fusion directly in the BEV space. The resulting multi-modal voxel features are merged by flattening the height and channel dimensions, followed by a 2D convolution to obtain BEV features. Although radar features are also fused in the voxel space, they are omitted from Figure 7 for clarity. SDMap fusion is performed separately in the BEV domain, where SDMap features are concatenated with the BEV features derived from other sensors. The mathematical formulation of the fusion pipeline, including voxelization, multi-modal concatenation, and SDMap integration, is detailed in Supplementary Section S.2 of the supplementary materials.
@@ 3.5 Auxiliary One-to-Many Set Prediction Loss Strategy
To improve training efficiency, an auxiliary one-to-many set prediction loss strategy [jia_detrs_2023, liao_maptrv2_2024] is employed. This study, to the best of our knowledge, is the first to perform ablation experiments on this strategy for road topology understanding. Results show that this approach significantly enhances the architecture’s ability to understand road topology.
The auxiliary one-to-many set prediction loss approach involves a smaller decoder and a larger decoder with shared weights. The smaller decoder uses the ground truth set directly, while the larger decoder employs a repeated ground truth set to increase the number of positive samples. During training, both decoders are utilized, while only the smaller decoder is employed during inference. The mathematical background of this loss strategy, including decoder sharing with masking, is provided in Supplementary Section S.3 of the supplementary materials.
@@ 4 Experimental Evaluation
This section provides a comprehensive analysis of the methodologies employed in this study. It commences with a detailed description of the datasets and metrics (Section 4.1). Then, the implementation details are provided in Section 4.2. Subsequently, Section 4.3 presents an extensive experimental evaluation, encompassing an analysis of instance mask formulation, comparisons of various attention mechanisms, an examination of auxiliary one-to-many set prediction loss, an efficiency analysis of the proposed Bezier deformable attention, and a study on multi-modal fusion involving camera, radar, lidar, and SDMap. The results are juxtaposed with state-of-the-art methods on both OpenLane-V1 and OpenLane-V2.
Supplementary material provides additional details on the various loss functions (Section S.4), the implementation specifics of the training setup, dataset preprocessing and the proposed architecture (Section S.5), and further experiments (Section S.6) including view transformation, backbone variations, number of epochs, number of encoder and decoder layers, and efficient multi-scale implementation ablations.
@@ 4.1.1 Datasets
Figure 8: Perspective-view (PV) and bird’s-eye-view (BEV) samples from the OpenLane-V2 dataset. (a) and (d) show centerline instances in PV and BEV domains, respectively, with each color representing a distinct instance. (b) and (c) illustrate centerlines with colors indicating topological relationships between centerlines and traffic elements in PV and BEV. (e) visualizes the topological relationships among different centerlines, where directed arrows indicate connectivity between centerlines.
The OpenLane-V2 dataset [wang_openlane-v2_2024] is a centerline detection and road topology understanding dataset and has two subsets: Subset-A and Subset-B. Subset-A is derived from the Argoverse 2 (AV2) [wilson_argoverse_2023] dataset, containing samples from six cities (e.g., Miami, Pittsburgh, Austin), collected using a seven-camera setup with the front camera positioned vertically. This subset contains 22,477 training samples, 4,806 validation samples, and 4,816 test samples, all with a resolution of $2048\times 1550$ pixels. Multi-modal fusion studies are conducted with lidar data and SDMap, and no radar data is available in Subset-A.
Subset-B is based on the NuScenes [caesar_nuscenes_2020] dataset and contains data collected from Boston and Singapore, using a six-camera setup. It includes 27,968 training, 6,019 validation, and 6,008 test samples, with an image resolution of $1600\times 900$ pixels. Subset-B has a higher proportion of night and rainy scenes compared to Subset-A, but lacks ground height information. Sensor fusion studies are implemented with both radar and lidar data, but no available SDMap information in this subset. The details of the image pre-processing pipeline for both subsets are provided in Supplementary Section S.5.4.
Figure 8 presents PV and BEV images from a random sample in Subset-A of the OpenLane-V2 dataset. Due to the vertical positioning of the camera, the front view in the PV domain is cropped. As shown in subfigures (a) and (d), centerline instances are visualized in PV and BEV domains, respectively, with each color representing a distinct instance. Subfigures (b) and (c) highlight the topological relationships between centerlines and traffic elements using color-coded representations in both PV and BEV views. Subfigure (e) illustrates the topological relationships among different centerlines, where directed arrows indicate connectivity between them. More ground truth examples are shown in Supplementary Figure S1 in Section S.9.
The OpenLane-V1 dataset [chen_persformer_2022] is a comprehensive benchmark for 3D lane detection, derived from the Waymo Open dataset. It consists of 1000 segments with 200K frames, captured under diverse conditions at $1920\times 1280$ resolution. This dataset provides diverse and challenging scenarios for evaluating lane detection algorithms, including a wide range of weather conditions, lighting variations, and road types.